[Torrey]
General updates from work in the lab so far today:
-The 1550 path is now well aligned to OFC2. The camera is very good in getting the alignment close, but minimizing the higher order modes is much more effective with a PD in transmission.
-I aligned a 1550 13 MHz bandwith PD to the flip up mirror TRANS path. This is quite annoying to do as the AOM has not been frequency matched to the 1550 light. There is almost no light transmitting.
-With the 1550 PD aligned I can minimize higher order modes resonant in the cavity. The alignment is now fine tuned and final, baring any minor touch ups as things get bumped. The configuration for the 1550 TRANS light can be see here: Down Up
-With a TRANS PD I can also match the AOM frequency such that the 775 and 1550 light are coresonant. The initial measurement can be seen here 3A193161-FFF8-4AB9-BD32-6A5FADD5ECE9.png. As seen the 1550 light is almost exactly midway between the 775 resonance peaks (at 200 MHz, this is the AOM optimal efficiency frequency). This means we need large AOM steps to get them coresonant.
-This new OFC2 cavity is using the new AOMs that arrived, they deflect the shifted beam way less. The one we borrowed from Ellie's lab could move +/- 1 MHz before completely losing all power going into the fiber. These can move +/- 15 MHz and still have some amount of power reaching the other side of the fiber. In fact, I can move a full 15 MHz, tune up the input coupler, move another 15 MHz, tune up, etc, very easily.** I've done this to make the 1550 and 775 light coresonant in OFC2. As seen here, the TRANS signal on the 1550 path and the zero point on the error signal on the 775 path are aligned. The frequency used to achieve this is 230.69 MHz.
** I'm not sure the progress on the curved mirror AOM experiment but, this may even prove the that unnecessary. I would still like to see how well the curved mirror works though.